# On energy bounds in asymptotically locally AdS spacetimes

### v1

@@ Appendix A Inverting the Dirac operator
$K=0$ is a “maximal gauge” that is sometimes used in the study of the initial value problem. However, it would have to be shown $(M,g)$ admits such a foliation and that the foliation is compatible with the coordinates chosen on $\mathcal{I}$ to get $f_{(0)0\alpha}=0$.
For the general case, without additional assumptions, consider the following index theory argument based on the analysis in [53, 39]. Let $\hat{\nabla}_{a}=D_{a}+\frac{\mathrm{i}}{2}\gamma_{a}$, i.e. don’t include the $\mathcal{A}_{a}$ term. Therefore, $\mathfrak{D}-\hat{\mathfrak{D}}=\gamma^{I}\mathcal{A}_{I}$. If $\gamma^{I}\mathcal{A}_{I}$ were a compact operator, then the index (dimension of kernel minus dimension of cokernel) of $\mathfrak{D}$ and $\hat{\mathfrak{D}}$ would coincide. In particular, if $\hat{\mathfrak{D}}$ was invertible, then it would follow that $\mathrm{index}(\mathfrak{D})=0$. Since $\mathfrak{D}$ has already been shown to have trivial kernel above, it would follow that $\mathfrak{D}$ is invertible.
Even though $\mathcal{A}_{a}$ is just a (sufficiently regular) matrix valued function, it may not be clear whether $\gamma^{I}\mathcal{A}_{I}$ is actually compact. This would be the case though if the underlying space, $\Sigma_{t}$, were itself compact - a fact leveraged in [53, 39].
[…]
Since $C_{c}^{\infty}$ is dense in $L^{2}$, for any $\Phi\in L^{2}$, $\exists$ a Cauchy sequence, $\{\phi_{m}\}_{m=0}^{\infty}\subseteq C_{c}^{\infty}$, such that $\lim_{m\to\infty}\phi_{m}=\Phi$ in $L^{2}$. For each $\phi_{m}$ choose an $r_{m}$ large enough so that $\mathrm{supp}(\phi_{m})\subseteq\mathrm{int}(\Sigma(r_{m}))$. Then, the choice of boundary conditions doesn’t matter and $\exists\psi_{m}$ such that $\mathfrak{D}(\psi_{m})=\phi_{m}$. The results in [50, 51] show $\mathfrak{D}$ has an “elliptic regularity” property on $\Sigma(r_{m})$ meaning $\psi_{m}$ will inherit the regularity of $\phi_{m}$.
Therefore, each $\psi_{m}$ is actually an element of $C_{c}^{\infty}$. Furthermore, $\{\psi_{m}\}_{m=0}^{\infty}$ is a Cauchy sequence using lemma A.4. By theorem A.6, the limit, $\lim_{m\to\infty}\psi_{m}=\psi\in\mathcal{H}$ satisfies
thereby proving $\mathfrak{D}$ is surjective.
@@
Suppose $\Phi$ is a spinor such that $\gamma^{I}\nabla_{I}\Phi\in L^{2}$ and $\Phi$ grows at most as $O(\mathrm{e}^{r/2})$. By theorem A.7 and the subsequent discussion, let $\Psi\in\mathcal{H}$ be the unique spinor such that $\gamma^{I}\nabla_{I}\Psi=\gamma^{I}\nabla_{I}\Phi$. Let $\mathcal{Z}=\Phi-\Psi$. Let $\{\psi_{m}\}_{m=0}^{\infty}\in C_{c}^{\infty}$ be a Cauchy sequence whose limit is $\Psi$ and let $\mathcal{Z}_{m}=\mathcal{Z}-\psi_{m}$. Then, for the functional, $Q$, in definition 2.5, $\lim_{m\to\infty}Q(\mathcal{Z}_{m})=Q(\mathcal{Z})$.

### v2

@@ Appendix A Inverting the Dirac operator
$K=0$ is a “maximal gauge” that is sometimes used in the study of the initial value problem [60]. However, it would have to be shown $(M,g)$ admits such a foliation and that the foliation is compatible with the coordinates chosen on $\mathcal{I}$ to get $f_{(0)0\alpha}=0$.
@@
Suppose $\Phi$ is a spinor such that $\gamma^{I}\nabla_{I}\Phi\in L^{2}$ and $\Phi$ grows at most as $O(\mathrm{e}^{r/2})$. By theorem A.7, let $\Psi\in\mathcal{H}$ be the unique spinor such that $\gamma^{I}\nabla_{I}\Psi=\gamma^{I}\nabla_{I}\Phi$. Let $\mathcal{Z}=\Phi-\Psi$. Let $\{\psi_{m}\}_{m=0}^{\infty}\in C_{c}^{\infty}$ be a Cauchy sequence whose limit is $\Psi$ and let $\mathcal{Z}_{m}=\mathcal{Z}-\psi_{m}$. Then, for the functional, $Q$, in definition 2.5, $\lim_{m\to\infty}Q(\mathcal{Z}_{m})=Q(\mathcal{Z})$.